OpenStudy (anonymous):

Choose the correct percent proportion used to solve "What is thirty percent of nineteen?

OpenStudy (imstuck):

Do you have choices from which to pick? It's \[\frac{ 30 }{ 100 }\times 19\]

jimthompson5910 (jim_thompson5910):

You can also write it as Part/Whole = Percent/100

jimthompson5910 (jim_thompson5910):

The "part" is unknown, so call it x The "whole" is 19 The "percent" is 30 giving us this x/19 = 30/100 in which you would solve for x to get your answer

OpenStudy (anonymous):

I normally just use \[\frac{ is }{ of }= \frac{ % }{ 100 }\] Just put in 30 in the blank space, put 19 in the of space. Then you multiply 30 by 19 and divide that number by 100 and you get your is